OpenStudy (anonymous):

2 over 3 next to x + 1 = 27

OpenStudy (hba):

\[\frac{ 2 }{ 3}(x+1)=27\] Right ?

OpenStudy (anonymous):

@hba yes(:

OpenStudy (hba):

1) multiply both sides by 3

OpenStudy (hba):

@sylinan Start working :)

OpenStudy (anonymous):

so we have 2/3(x+1)=27 . We multiply the opposite of 2/3 to both sides, which is 3/2, so the fraction can cancel out and that means we have- \[3/2∗2/3(x+1)=27∗3/2 \] \[x+1=27∗3/2 \] Now we multiply 27/1*3/2 and that gives us 81/2 and 81/2=40.5 \[x+1=40.5 \] Now we finally then subtract both sides by 1 which gives us \[x=39.5\]

OpenStudy (anonymous):

@Loujoelou thank you(: